OpenStudy (anonymous):

A total of 276 tickets were sold for the school play. They were either adult tickets or student tickets. The number of student tickets sold was two times the number of adult tickets sold. How many adult tickets were sold?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Hmmm didn't we do this one? Well set it up like the others... a + s = 276 s = 2a a + (2a) = 276 3a = 276 a = ?

OpenStudy (anonymous):

92

OpenStudy (johnweldon1993):

Perfect!

OpenStudy (anonymous):

92 adult tickets were sold?

OpenStudy (johnweldon1993):

Yes...because if 92 adults were sold...that means that twice that (184) students were sold....so 92 + 184 = 276 which is what they sold